Solve the system of equations using the addition method
-4x+6y=-6
5x-3y=36
I completely understand how to work this problem until the step to solve for y, which should be a fraction but I don't know how to correctly solve it. Any help will be much appreciated!!

Answer by Alan3354(69443)

-4x+6y=-6
5x-3y=36
--------
-2x+3y=-3 Eqn 1 x 1/2
5x-3y=36
--------- Add
3x = 33
x = 11
-------
5x-3y=36
5*11 - 3y = 36
55 - 3y = 36
-3y = -19
y = 19/3